OpenStudy (anonymous):

please help!!!!!!!!! 2A+3B-4C=7 A-5B+2C=2 -7A-3B+5C=-20 and my prof. answer is (1,-1,-2). I could not get the same answer with him. I tested the answer the first 2 equaltions work out, but the third seems not. Please I really need help!!!!!!!!

OpenStudy (anonymous):

I dont know that method. Do you mind helping me step by step:)

OpenStudy (anonymous):

I got A=141/57 B=-9/57 the numbers are not pretty so I stop

OpenStudy (anonymous):

Oh thank you I will watch it:)

OpenStudy (anonymous):

But do you think my Prof. 's answer is not correct???

hartnn (hartnn):

-7A+3B+5C=-20 that is the 3rd equation

OpenStudy (anonymous):

@hartnn yes

hartnn (hartnn):

u mentioned it incorrectly in question....

hartnn (hartnn):

maybe u used that incorrect equation to solve...

OpenStudy (anonymous):

no the 3rd equation is -7A-3B+5c=-20

hartnn (hartnn):

how are u so sure ?

hartnn (hartnn):

2A+3B-4C=7 A-5B+2C=2 -7A+3B+5C=-20 -------------> (1,-1,-2)

OpenStudy (anonymous):

3rd: -7A -3B+5C=-20 not -7A+3B+5C=-20

OpenStudy (anonymous):

@ktnguyen1 if you have the third equation as you wrote it at first (it's written wrong) then the answer you got for A and B (the ugly numbers) are right. If you use the corrected third equation, as @hartnn said, then the professor's answer is the right one.

hartnn (hartnn):

2A+3B-4C=7 A-5B+2C=2 -7A-3B+5C=-20 ---------->(47/19,-3/19,-12/19)

OpenStudy (anonymous):

yes, I think his hand written was wrong!@hartnn, @nincompoop: are correct

OpenStudy (anonymous):

Thank so much guys!!!